OpenStudy (jkbo):

Determine if the conditional is true or false: If two numbers are odd, then their sum is even. (1 point) true false

Parth (parthkohli):

3 + 3 = 6 <--- Even 1 + 5 = 6 <---- Even @GOODMAN 3 + 3 = 9? o.O

OpenStudy (jkbo):

so its true?

Parth (parthkohli):

Yep, it is :)

OpenStudy (anonymous):

Wow..lol. I multiplied, my bad. lol

OpenStudy (jkbo):

thank u guys

Parth (parthkohli):

yw :)